### v1

# Introducing Milabench: Benchmarking Accelerators for AI

Abstract
AI workloads, particularly those driven by deep learning, are introducing novel usage patterns to high-performance computing (HPC) systems that are not comprehensively captured by standard HPC benchmarks. As one of the largest academic research centers dedicated to deep learning, Mila identified the need to develop a custom benchmarking suite to address the diverse requirements of its community, which consists of over 1,000 researchers. This report introduces Milabench, the resulting benchmarking suite. Its design was informed by an extensive literature review encompassing 867 papers, as well as surveys conducted with Mila researchers. This rigorous process led to the selection of 26 primary benchmarks tailored for procurement evaluations, alongside 16 optional benchmarks for in-depth analysis. We detail the design methodology, the structure of the benchmarking suite, and provide performance evaluations using GPUs from NVIDIA, AMD, and Intel. The Milabench suite is open source and can be accessed at github.com/mila-iqia/milabench.
1 Introduction
Recent breakthroughs in AI have been driven by a combination of factors, including increased computing power, the availability of large datasets, and advances in deep learning research. Graphics Processing Units (GPUs) have become a cornerstone of modern AI pipelines, with neural network architectures having co-evolved to leverage the high parallelization capacity of GPUs. Given that hardware and energy costs in HPC systems optimized for AI workloads are predominantly driven by GPUs, selecting the right GPU is critical for achieving optimal system performance within budget constraints.
AI workloads are inherently diverse, presenting unique challenges that are difficult to replicate accurately using simple kernels or synthetic benchmarks. To address this, Mila began developing an internal benchmarking suite in 2019 to support its procurement process for an internal cluster dedicated to its AI researchers. Over time, this benchmarking suite, Milabench, has matured into a robust solution, which we present in this report.
The latest version of Milabench draws upon a comprehensive literature review of 867 papers published by Mila researchers in 2023. It encompasses a broad spectrum of AI research domains, model architectures, and scaling pipelines, including 26 main benchmarks for procurement evaluations and 16 additional benchmarks for deeper analysis.
Milabench is designed and implemented based on 3 principal objectives:
  1. 1.
    Simple to use: To streamline the procurement process.
    Fast
    Short test durations enable rapid experimentations.
    Modular
    A uniform interface across benchmarks simplifies both execution and contributions.
  2. 2.
    Representative: To ensure the benchmarks reflect real-world usage.
    Diverse
    Covers a wide range of research topics and models studied at Mila.
    End to end
    Addresses bottlenecks beyond GPU performance, reflecting the full pipeline.
    Closed
    Avoids vendor-specific optimizations that researchers are unlikely to use.
    Dependencies
    Incorporates libraries commonly used by researchers to validate compatibility.
  3. 3.
    Unbiased: To promote diversity in hardware and software solutions.
    Thorough Testing
    Benchmarks are validated across multiple hardware vendors.
    Generic Pipelines
    Favors widely adopted alternatives over vendor-contributed pipelines to ensure generality.
We will describe in Section 2 the design strategies used to select the benchmarks in the latest version of Milabench, along with a full list of the benchmarks included. A brief overview of the implementation of the suite will be covered in Section 3, before showcasing results in Section 4. We will conclude by providing a comparison with other benchmarks in Section 5, and discuss future directions in Section 6.
2 Suite Design
The benchmarks included in Milabench are carefully selected and balanced to reflect the breadth of research conducted at Mila. With a community of over 1,000 researchers, professors, and graduate students, the institute spans a diverse range of research topics. The selection process is informed by surveys and an extensive literature review of 867 papers published by Mila members in 2023.
This section begins with a detailed presentation of the literature review (2.1), followed by a discussion of the survey results (2.2). Finally, we outline the selection strategy derived from these analyses (2.3).
2.1 Literature Review
The primary source for benchmark selection is the comprehensive literature review. A total of 867 papers published by Mila researchers in 2023 were collected using Paperoni, an open-source utility developed at Mila, and converted to raw text using pdftotext [31]. Annotation of the papers was performed using GPT-4o, extracting key information such as research domains, model descriptions, and dataset details. An example of a response from GPT-4o is provided in Appendix B.
Evaluation
To validate the accuracy of GPT-4o and assess the utility of the resulting statistics, we manually annotated 110 papers published prior to 2024. To ensure a broad distribution of research topics, a pool of 750 papers published between 2019 and 2024 was initially categorized using a naive approach: literal keyword matching of 11 predefined domains within their abstracts [^1] Audio, Computer Vision, Generative Models, Graph Neural Network, Medical, Multimodal, Natural Language Processing, Neuroscience, Recommendation System, Reinforcement Learning, Robotics,. From each category, 10 papers were randomly sampled, resulting in a balanced set of 110 papers. These 110 papers were processed twice by GPT-4o, yielding 220 responses. Both responses for each paper, along with the corresponding PDF, were manually reviewed, corrected, and merged to create a high-quality annotation set.
To assist in validating GPT-4o’s predictions, the JSON template provided to the model included fields for justifications and quotes supporting each value. This strategy was applied to both research domains and model names to enhance transparency and interpretability.
The validation set was subsequently used to evaluate GPT-4o’s performance via a Multi-Label Confusion Matrix (MLCM) [11], focusing on the accuracy of predictions for research domains and model types.
Predictions often exhibited significant variability across papers, with frequent overlaps between domains and model architectures. Additionally, many papers described numerous models, which GPT-4o rarely captured comprehensively. To address these challenges, we grouped models into broader architecture families rather than requiring precise identification. A hierarchical taxonomy of domains and models was manually created to aggregate statistics at a higher level of abstraction. For instance, specific models like Llama were mapped to broader categories such as Transformer. Table 1 presents quality measures for GPT-4o’s predictions compared to manual annotations, aggregated using this hierarchy. In this evaluation, predictions were mapped to the hierarchy’s top level. For a given paper, a top-level category was assigned a count of 1 if there was at least one match and 0 otherwise. This aggregation strategy provided a more comprehensive overview and improved both recall and precision by over 20%. The hierarchy included numerous top-level categories beyond those shown in Table 1. Less descriptive categories (e.g., Deep Learning Optimization or Neural Network) and those with sparse occurrences were consolidated into the category Others.
Results
We observe recall rates ranging from 84% to 98% and precisions ranging from 86% to 100% for domains, with averages of 94% for both metrics. For model architectures, recalls ranged from 50% to 100% and precisions ranged from 84% to 100%, with an average of 75% and 96% respectively.
Table 1: MLCM tables for the domains and model with other domains or model architectures identified by GPT-4o. Additional column and row NPL and NTL stands for No Predicted Label and No True Label respectively.
Domains Model Types
Predicted Predicted
Computer Vision Graphs Natural Language P. Reinforcement L. Others NPL CNN Diffusion model Gflow nets GNN MLP RNN Transformer N/A NPL
True Labels Computer Vision 15 0 0 0 1 2 CNN 23 0 0 0 0 0 0 1 6
Graphs 0 6 0 0 2 2 Diffusion model 0 4 0 0 0 0 0 0 1
Natural Language P. 0 0 17 0 0 4 Gflow nets 0 0 5 0 0 0 0 0 0
Reinforcement L. 0 0 0 15 3 3 GNN 0 0 0 10 0 0 0 0 4
Others 1 1 1 0 92 7 MLP 0 0 0 0 2 0 0 0 1
NTL 0 0 0 0 0 0 RNN 0 0 0 0 0 3 0 0 3
Transformer 0 0 0 0 0 0 38 2 6
N/A 0 0 0 1 0 0 0 48 18
NTL 1 0 0 0 0 0 1 6 0
Precision (%) 94 86 94 100 94 96 100 100 91 100 100 97 84
Recall (%) 98 97 97 96 84 77 80 100 71 67 50 83 72
These results served as targets to balance the benchmarks in Milabench, ensuring they represent the relative proportions of research domains and model architectures used by Mila researchers. Figure 3 illustrates these proportions compared to the benchmarks included in the suite.
One notable finding from the literature review is the significant proportion of research focused on or involving Reinforcement Learning (RL), a domain previously absent from earlier versions of Milabench. While Computer Vision and Natural Language Processing (NLP) were expected to dominate as the most prominent research domains, the results reveal that NLP and RL hold similar levels of importance at Mila. An additional insight is the nuanced role of Computer Vision in research. Only about 19% of papers were explicitly categorized as focusing on CV by GPT-4o. However, an additional 18% of papers used vision datasets without being explicitly categorized under CV, often being labeled as Deep Learning Neural Network or Deep Learning Optimization. This brings the total proportion of papers involving CV tasks to 37%. Of these, roughly half focus on CV-specific studies, while the other half employ CV tasks as a general experimental platform. A similar pattern, though less pronounced, was observed for NLP. Approximately 4.4% of papers utilized NLP datasets without being categorized under NLP. Conversely, this phenomenon is negligible in the case of RL, where papers generally align directly with their core research domain.
The proportion targets for research domains used in the design of benchmarks, as shown in Figure 3 (blue bars), are derived from a combination of domain categorizations and dataset domain categorizations. In contrast, the proportion targets for model architectures are based directly on the statistics obtained from GPT-4o.
2.2 Surveys
Surveys were conducted to address gaps in the literature review and to account for emerging trends. These surveys primarily aimed to gather statistics on the most commonly used libraries and to identify potential benchmark pipeline candidates.
Libraries
Only A small fraction of papers explicitly cite or mention the libraries used for the experiments presented. As a result, we relied on internal surveys to identify the most widely used libraries for inclusion in Milabench.
PyTorch is by far the most commonly used library, with 96% of Mila researchers adopting it. In contrast, TensorFlow is used by only 5% of researchers, all of whom also use PyTorch. The adoption of JAX has been steadily increasing, with 26% of researchers now using it. Among these, 2.6% use JAX exclusively, while the remaining 23.4% use both JAX and PyTorch. Given these usage patterns, we decided to include tests using only PyTorch and JAX. JAX is primarily utilized for reinforcement learning research, a trend that is reflected in the suite.
Both NLP and Computer Vision researchers have increasingly adopted Hugging Face, with approximately 33% of researchers using it. Additionally, the generic framework PyTorch Lightning is used by 26% of Mila researchers.
Benchmark pipeline candidates
Professors were surveyed to suggest new pipelines for Milabench. However, the responses were highly biased, underscoring the need for a more reliable method to gather statistics on the domains to be covered. A comparison of survey results with our literature review revealed notable discrepancies. For instance, 60% of the suggested pipelines focused on Computer Vision, while only 37% of papers in the literature review were dedicated to this domain. Similarly, 15% of suggested pipelines were for NLP, compared to 26% in the literature, 20% for RL versus 26%, and 5% for Graphs versus 15%.
Given these biases, we relied on the expertise of the professors to suggest more pipelines than necessary but ultimately used the literature review to make the final selection.
A total of 25 pipelines were suggested, from which 10 were chosen. We reused 6 pipelines from the previous version of Milabench and added 3 additional pipelines to address obvious gaps.
2.3 Selection of benchmarks
Coverage targets for the benchmarking suite were defined across several dimensions, including research domains, model architectures, model sizes, training parallelization methods, and libraries. These targets are not mutually exclusive, meaning that some benchmarks can cover multiple areas simultaneously, allowing for broader coverage with fewer tests.
The pipelines were selected based on the popularity of models and datasets, as well as the availability of mature open-source implementations that are well-supported by the research community. We rely on full implementations from third-party libraries and make only minimal modifications to integrate them into Milabench. Additionally, we avoid using models or frameworks contributed by vendors to minimize potential biases in favor of those vendors.
The main benchmarking suite consists of 26 benchmarks, encompassing 19 different model architectures. For certain model architectures, there are multiple variations in execution pipelines to reflect the diverse usage patterns of researchers. For example, benchmarks using Llama include variations for pre-training with different parallelization budgets, fine-tuning with LoRA also including different parallelization budgets, and an inference tasks, totaling 6 distinct benchmarks. The selection of main benchmarks and their corresponding targets is illustrated in Figure 3. A detailed list of the main benchmarks can be found in Table 2.
In addition to the 26 main benchmarks, 16 optional benchmarks are available for further analysis. These include variations in precision formats (fp16, bf16, tf32, and fp32) for basic matrix multiplication, transformer training (BERT), and CNN training (ConvNext). The benchmark resnet50 is sensitive to I/O and CPU performance, so we included an additional version without I/O to measure the impact of potential bottlenecks caused by low bandwidth or insufficient CPU resources.
Some main benchmarks are only included as multi-GPU tasks (e.g., dinov2-giant-gpus) because training these models on a single GPU would be too slow for practical use. For these, we provide single-GPU versions as optional benchmarks, allowing for scaling measurements from single-GPU to multi-GPU setups within a single node, and from single-node to multi-node configurations (e.g., diffusion-nodes). While rlhf-single would also be more efficiently trained using multiple GPUs, we include the single-GPU version as the main benchmark and the multi-GPU version as an optional benchmark to avoid further biasing the suite toward single-node training.
A detailed list of the optional benchmarks is provided in Table 3. The optional benchmarks are not represented in Figure 3, as they do not contribute to achieving the coverage targets.
3 Implementation
Milabench is composed of four main components, milabench, benchmate, voir and torchcompat.
  • milabench: manage the benchmarking process, from its configuration to its execution, including the installation of dependencies and the downloads of datasets.
  • benchmate: common utilities that can be reused between benchmarks, such as custom TimedIterators or custom monitors to log metrics.
  • voir: instrumentation tool to retrieve and log metrics. It includes GPU monitors for different vendors and custom multiprocessing tools to help log metrics.
  • torchcompat: compatibility layer to smooth out differences between vendors. It implements a superset of torch.cuda to enable all vendors to run seamlessly with no code change.
Each benchmark in Milabench is implemented as an independent "package." These packages are designed to be isolated from one another, allowing each to install its dependencies within its own virtual environment to prevent dependency conflicts.
However, many benchmarks share common dependencies, and to optimize resource usage, we have consolidated these shared dependencies into a single virtual environment. This approach not only reduces the overall size of Milabench but also simplifies execution. In practice, we strive to use a single environment across benchmarks, which simplifies result analysis and helps keep Milabench installation lightweight.
Milabench includes helper commands to manage dependency versions easily, allowing them to be pinned to the latest compatible releases.
The execution of Milabench is organized into three steps. The first two steps comprise an initial setup process, which can be skipped after they have been successfully run once.
  • install: Create virtual environment for each benchmarks and install their dependencies
  • prepare: Download checkpoints and datasets that will be required to run the benchmark suite
  • run: Execute the benchmark suite
  • report [Optional]: Generate an aggregate report of all the runs
The benchmark suite is fully defined by a YAML configuration file, which specifies the benchmarks to be executed and their execution parameters. Users can create customized benchmark suites by selecting from the pool of available benchmarks.
Additionally, we provide a set of Docker images to simplify benchmark execution. These images come with pre-installed dependencies to ensure better reproducibility of results.
3.1 Benchmarking Methodology
To provide metrics that closely reflect real-world scenarios, the benchmarks in Milabench are designed to mirror the experimental pipelines commonly used by researchers. This approach requires not only selecting the most popular models and libraries but also ensuring that the code closely follows typical research workflows, with minimal adaptations for vendor support and without introducing additional optimizations. This design philosophy also simplifies the performance measurement process.
By relying on third-party libraries that are well-supported by the community, we focus on measuring system throughput rather than optimizing for convergence time in training objectives. As a result, the 42 benchmarks in Milabench can be executed in under 2 hours222assuming datasets and checkpoints have been pre-downloaded on the latest available GPUs.
Each benchmark is implemented as a short execution to gather 30 to 60 measurements of processing time for units of work. The nature of these units of work varies depending on the benchmark and can be defined in terms of the number of samples, tokens, nodes, or steps within episodes.
Where possible, datasets are generated randomly to minimize reliance on external data sources. However, in some cases, subsets of standard datasets are used when the characteristics of the data affect processing time. For example, in LLM tasks, pure random noise would result in the LLM consistently producing maximum-length sequences. The randomly generated datasets are sized to match popular real-world datasets; for instance, FakeImageNet consists of images sized at (384, 384), based on the average image sizes in ImageNet [4], and FakeVideos consists of 30 FPS, 10-second videos at (640, 480), modeled after the Kinetics dataset [13].
This approach is also applied to model initialization, particularly in inference mode. Using a pre-trained model ensures performance metrics are aligned with real-world scenarios, as using an untrained model would significantly skew results.
Pipelines that typically incorporate online data augmentation are implemented as such, with the augmentation process affecting the overall performance measurement. Otherwise, data preprocessing is handled during the preparation phase and is not included in the performance evaluation.
Tasks may be executed on three different scales: 1) single-gpu, 2) single-node multi-gpu or 3) multi-node multi-gpu. In all cases, all available GPUs are used to measure full-node performance. For single-GPU tasks, the benchmark is launched separately on each GPU, and performance is averaged across GPUs. This approach helps identify issues such as insufficient CPU resources, which may hinder GPU performance. For multi-GPU tasks, whether on a single node or across multiple nodes, performance is measured across the GPUs without any normalization.
Scores are aggregated across all main benchmarks using a weighted geometric mean, implemented via the logsumexp function for improved numerical stability. Benchmarks performances are multiplied by their success rate, then increased by 1 to avoid $\log 0$. This gives the following equation:
$\exp\left(\frac{\sum_{i}w_{i}\log{(p_{i}*s_{i}+1)}}{\sum w_{i}}\right)$ (1)
Where $i$ is the benchmarks index, $w_{i}$ is a weight, $p_{i}$ is a performance measure and $s_{i}$ is a success rate.
4 Results from Milabench
We executed Milabench across four different node pairs, each equipped with distinct GPU models: Nvidia A100 SXM4 80GB, Nvidia H100 SXM5 80GB, AMD MI300X 192GB, and Intel Gaudi2 96GB. The performance results from the A100 were used as the baseline for comparison with the H100, MI300X, and Gaudi2 GPUs. For further details on the node specifications, please refer to Table 6.
4.1 Optional Benchmarks
Before delving into the main results from Milabench, we first examine the FLOP performance for each vendor’s GPUs (Figure 5). The data reveals distinct design priorities across vendors. Both the Gaudi2 and MI300X emphasize float32 performance, whereas the H100 focuses on low-precision operations, particularly with TF32.
In terms of low-precision performance, the H100 and MI300X appear to be on par, while the Gaudi2 lags behind  [^3] It is worth noting that the Gaudi2 was released in May 2022, while the H100 and MI300X were released in March 2023 and December 2023, respectively..
In the synthetic FLOP benchmark, the standout feature of the H100 is its performance with TF32. While the other vendors support TF32, they do not show a significant performance improvement over the A100. The official MI300X specification mentions TF32, but we did not observe any notable gains compared to single-precision performance. This could be due to a lack of support in PyTorch. The Gaudi2, on the other hand, may default to using TF32 in PyTorch without an option to disable it, which could obscure its performance characteristics.
Despite the strong performance of MI300X and Gaudi2 observed in the synthetic FLOP benchmark, the results do not fully translate into real-world workloads, such as BERT and ConvNext training (Figure 5). The H100 outperforms both the MI300X and Gaudi2 in low-precision settings, even though it was comparable to the MI300X in the synthetic low-precision (FP16) benchmark. Moreover, in single-precision tasks, the H100 demonstrates competitive performance with the MI300X and Gaudi2, despite the latter two GPUs achieving significantly higher FLOP counts in the synthetic benchmark.
The full results for optional benchmarks are presented in Table 5.
Figure 1: Performance comparison between different data types [^5] Note that Gaudi2 might be using TF32 on the fp32 benchmark.
4.2 Main benchmarks
Finally, as seen in the result overview in Figure 7, the trends observed for BERT and ConvNext generally extend across other domains as well. Despite the MI300X matching the H100 in terms of low-precision FLOPS and outperforming it in single-precision performance, the MI300X is consistently outperformed in all benchmarks.
We attribute this performance disparity between FLOP counts and real-world performance to differences in software stack maturity. CUDA has been the de facto standard for AI and machine learning workloads for over a decade. During this time, its compute kernels have been continually tuned and developed alongside AI advancements, and it remains tightly integrated with the most popular machine learning libraries. As the default platform, CUDA naturally benefits from extensive testing and optimization. In contrast, newer platforms need to invest considerable effort to catch up to years of innovation and optimization inherent in CUDA.
AMD and Intel have taken different approaches to address this challenge. AMD appears to have chosen a broad strategy, emulating CUDA and focusing on achieving feature parity. This approach is evident in the wide test coverage, with only one benchmark unsupported. Intel, on the other hand, has opted for a more targeted approach, focusing on high-demand applications and developing highly optimized, customized implementations. Milabench’s design favors the first approach, as it aims to faithfully represent common experimental pipelines, avoiding vendor-specific re-implementations. If a benchmark is supported in Milabench, it indicates that there is a low barrier to entry for working with the corresponding GPUs in that research domain.
Figure 2: Results Overview  [^7] Note that Gaudi2 did not support any of the graphs benchmarks
Natural Language Processing
NLP has gained significant popularity in recent years, with Large Language Models (LLMs) becoming a primary focus for all vendors. While the performance gap across GPUs in NLP is smaller compared to other domains, there is still noticeable variation. All GPUs perform well with full Llama pre-training, but fine-tuning with LoRA presents challenges for the MI300X and Gaudi2. We anticipate that major improvements will be seen in this area in future releases.
Computer Vision
Computer Vision is the most supported domain, with all vendors performing well on vision models. Older models like ResNet50 are particularly well-optimized. Newer or less popular models, however, perform better on the H100.
Reinforcement Learning
Historically less compute-intensive than other domains, RL is now leveraging GPUs for environment execution using JAX, introducing new usage patterns for GPU benchmarks. There is yet major improvements to come in this field from every vendors. H100’s averaged result is boosted mainly by the 2 tasks recursiongfn and rlhf-single that involves transformers, but makes its smallest gains of all benchmarks over A100 on the more canonical tasks.
Graphs
The Graph domain has been rapidly evolving, making support more challenging. Graphs typically rely on sparse operations, which are not supported by all vendors. Moreover, many custom CUDA kernels are built during installation, which complicates vendor support. Despite these challenges, the MI300X successfully ran all benchmarks, while Gaudi2 lacked support for compiling custom CUDA kernels into a compatible format.
The full results for main benchmarks are presented in Table 4.
The AI field is evolving quickly, and the results presented here are likely to become outdated soon. New PyTorch versions often bring performance improvements that range from small 5% gains to more than 100% on specific benchmarks. These performance gains can be attributed to algorithmic improvements that apply to all vendors but can also be attributed to improved integrations with some vendor stacks. Based on the discrepancies observed between synthetic and real-world benchmarks (as discussed in Section 4.1), we expect future software updates to significantly impact GPU performance. This discrepancy seems to stem more from differences in software stack support than hardware limitations. In short, stay tuned for result updates.
5 Related Work
High-Performance Computing (HPC) has a long-standing tradition of benchmarking, with tools like LINPACK [6] still being used to determine the TOP500 list of the world’s most powerful computer systems. Although LINPACK offers valuable insights into the raw performance of a system, it does not always capture the performance characteristics observed in real-world applications.
In recent years, HPC has more widely adopted procurement practices relying strongly on application-based benchmarks in addition to pure kernel benchmarks. This transition aims to provide a more accurate assessment of system performance, aligned with the anticipated usage patterns of new systems. For example, the Jülich Supercomputer Center (JSC) has developed a benchmark suite consisting of 16 application benchmarks and 7 synthetic benchmarks [10]. Of these, 3 are AI-focused. These benchmarks include: MMoCLIP, A multi-modal text-image task using a ViT-L-14 model, trained across 8 nodes for a total of 32 GPUs; Megatron-LM, A 175 billion-parameter language model trained on a language modeling task, trained across 96 nodes for a total of 384 GPUs; and ResNet, A ResNet50 model trained for image classification, executed on 10 nodes for a total of 40 GPUs. MMoCLIP and Megatron-LM are excellent candidates to benchmark large-scale training. ResNet on the other hand uses a model that is arguably too small to benefit from large scale training on latest GPUs. The benchmark is missing important domains of AI such as RL and Graph learning. This may be intentional however as the benchmark is tailored for large scale training whereas RL and Graph learning pipelines are more generally executed on a single GPU or a single node. While MMoCLIP and Megatron-LM are well-suited to benchmark large-scale AI training, the ResNet benchmark is based on a model that may be too small to fully exploit the capabilities of modern GPUs in large-scale training scenarios. Moreover, the benchmark suite does not cover critical AI domains such as reinforcement learning (RL) and graph learning. This omission is likely intentional however, as these domains are typically executed on a single GPU or node, whereas the suite is specifically tailored for large-scale training environments.
The benchmark requirements for general HPC system procurement must cover a broad spectrum of research areas, which makes it challenging to allocate more than a few tests specifically for AI. In contrast, for AI-focused benchmarking, MLPerf [26], developed and supported by MLCommons, serves as a key reference. Initially focused on training benchmarks, MLPerf has since expanded to include benchmarks for inference [27] and for HPC systems [7]. Rather than prescribing specific implementations, MLPerf provides guidelines and rules for implementing pipelines, thereby allowing vendors the flexibility to introduce optimizations and re-implementations. This approach fosters innovation by allowing significant modifications while still ensuring the model quality meets the target standards. However, this flexibility comes at the cost of potentially poor reproducibility of performance results, the vendor-specific optimizations being trapped in vendor implementations rather than contributed to popular libraries supported by the community of researchers. In contrast, Milabench emphasizes the use of precise, standardized implementations to ensure robust reproducibility and greater generalizability of performance across similar pipelines. Another notable distinction between MLPerf and Milabench is the reporting methodology. MLPerf allows vendors to submit results for individual benchmarks, without providing a unified framework for aggregating these results into an overall performance measure. While this fragmented approach may benefit vendors who wish to highlight specific strengths, it creates a disjointed landscape of results, making it difficult to draw general comparisons across different vendors. Milabench, on the other hand, adopts a more holistic approach, striving for comprehensive coverage across benchmarks. This ensures that researchers at Mila are not left behind when acquiring hardware that may not fully support their research tools. Consequently, Milabench aims to acquire complete sets of results, where any missing benchmark negatively impacts the overall score, ensuring that the system’s performance is evaluated in a manner that is both comprehensive and consistent.
6 Future Directions for Milabench
We aim to update Milabench annually to ensure that its benchmarking pipelines reflect current research advancements and incorporate support for newer hardware as it becomes available. However, obtaining hardware from multiple vendors for testing purposes remains a significant challenge, which has limited our ability to maintain continuous integration across a diverse set of hardware platforms. To address this, we are actively working to expand support for a broader range of vendors and to develop diagnostic tools that can assist in identifying potential issues and optimization opportunities for these systems.
Currently, the literature review process that underpins Milabench still requires substantial manual intervention, particularly in curating domains and model architectures into a structured hierarchical taxonomy. To scale the review process and enable it to handle a larger corpus of publications, we are exploring automation techniques that would allow the literature review to be updated annually, streamlining the incorporation of new research.
Additionally, the current classification of model architectures in Milabench does not fully account for the diversity of operations within individual models. For example, many models include dense layers at the output stage, which are structurally similar to Multi-Layer Perceptrons (MLPs). Autoencoders, too, are often integrated with other models or serve as components within larger architectures. We believe that rethinking the categorization of model architectures to focus on individual components or "architecture parts" will provide a more accurate representation of the computational patterns prevalent in AI pipelines.
Furthermore, there has been limited attention given to variations in input sizes beyond the selection of popular dataset formats. Future benchmarking targets should incorporate not only the design of input/output sizes but also include optional benchmarks that measure the impact of varying input sizes across different GPUs. This would help evaluate the scalability and performance sensitivity of models to input size changes, which is crucial for optimizing AI workloads on modern hardware.
Finally, as energy consumption becomes an increasingly important concern with the proliferation of densely packed HPC systems, we recognize the need to incorporate energy efficiency into the global performance metric of Milabench. Future versions of the benchmark suite should account for energy costs in addition to raw performance, providing a more comprehensive assessment of system efficiency and sustainability in high-performance AI and HPC environments.
7 Conclusion
Milabench offers a comprehensive solution that allows vendors to quickly test and experiment with different configurations across a wide range of models. At the same time, it gives customers the flexibility to design custom benchmarks tailored to their specific needs, enabling them to choose the hardware that best aligns with their requirements.
Beyond providing insights into hardware performance, Milabench also evaluates the maturity of the vendors’ software stacks. This dual focus ensures that the needs of research institutions and compute centers are thoroughly addressed. Additionally, it provides vendors with valuable feedback on how to improve their software stacks, highlighting areas with weak coverage, potential blind spots, and suboptimal compute kernels. This helps vendors optimize both their hardware and software solutions more effectively.
Appendix A Hardware & Configuration
{adjustwidth}
[…]
Table 6: Node specifications used for the results
Note that due to timeline issues, not all vendors could be ran with the same version of PyTorch. We use PyTorch 2.4 version or the latest available at the time. PyTorch version can have a major impact on performance.
Appendix B GPT-4o Response Example
"title": {
  "value": "UNMASKING THE LOTTERY TICKET HYPOTHESIS: WHATS ENCODED IN A
  WINNING TICKETS MASK?",
  "justification": "This is the exact title of the paper provided in the
  given text.",
  "quote": "UNMASKING THE LOTTERY TICKET HYPOTHESIS: WHATS ENCODED IN A
  WINNING TICKETS MASK?"
"primary_research_field": {
  "name": {
    "value": "Model Pruning",
    "justification": "The primary focus of the paper is on understanding
    iterative magnitude pruning (IMP) and its effectiveness in identifying
    sparse subnetworks that can be trained to achieve the same accuracy as
    the full network.",
    "quote": "A state of the art algorithm for doing so is iterative
    magnitude pruning (IMP) (Frankle et al., 2020a)."
"sub_research_fields": [
    "name": {
      "value": "Neural Network Training",
      "justification": "The research extensively analyzes how training
      procedures, such as gradient descent, interact with pruned networks
      and the robustness of these methods.",
      "quote": "Second, how does SGD starting from the masked rewind point
      extract and use this information?"
"models": [
    "name": {
      "value": "ResNet-20",
      "justification": "ResNet-20 is one of the network architectures used
      for empirical investigations in this paper.",
      "quote": "We do so through extensive empirical investigations on
      [...] modern network architectures (ResNet-20, ResNet-18, and
      ResNet-50)."
    "name": {
      "value": "ResNet-18",
      "justification": "ResNet-18 is one of the network architectures used
      for empirical investigations in this paper.",
      "quote": "We do so through extensive empirical investigations on
      [...] modern network architectures (ResNet-20, ResNet-18, and
      ResNet-50)."
    "name": {
      "value": "ResNet-50",
      "justification": "ResNet-50 is one of the network architectures used
      for empirical investigations in this paper.",
      "quote": "We do so through extensive empirical investigations on
      [...] modern network architectures (ResNet-20, ResNet-18, and
      ResNet-50)."
Appendix C Methodology
Milabench use hardware event as timer to ensure we time the execution of the kernels in its entirety. To avoid unnecessary sync, the events are appended during training and only logged at the end of the epochs.
To simplify the addition of new benchmarks, an iterator wrapper TimedIterator is provided, which encapsulate the necessary logic.
for i in range(epoch):
events = []
# Creation of the iterator from the dataloader is time consuming
# it would get amortized across many batch during real training
batch_iter = iter(loader)
total_obs = 0
# Avoid sync in the batch loop
start = Event()
start.record()
for batch in batch_iter:
pred = model(batch)
loss = fn(pred, target)
end = Event() # <-+ Limited
end.record() # | overhead
obs = (start, end, len(batch), loss.detach()) # | happens
events.append(obs) # | during
if len(events) + total_obs >= 60: # | kernel
break # | runs
start = end # <-+
# Force sync at the end of the epoch # <-+
for start, end, bs, loss in events: # | Timer is off
end.wait() # | logging does not
log(loss=loss.item()) # | impact
log(rate=bs / (end - start)) # | perf
total_obs += len(events) # | measures
if total_obs >= 60: # |
raise StopProgram() # <-+
Additionally, most of milabench benchmarks are research repository, to avoid having to fork or modify the original repository milabench use libraries to hot patch the code where necessary to use milabench tools to record and measure performance. This allows milabench to dynamically wrap dataloaders with its own TimedIterator without modifying the original code.
This simplifies the update of benchmarks (as we can simply pull the changes from the original repo) and ensures the benchmark is as close as possible to the real world use case.

### v2

@@
 Given these biases, we relied on the expertise of the professors to suggest more pipelines than necessary but ultimately used the literature review to make the final selection.
-A total of 25 pipelines were suggested, from which 10 were chosen. We reused 6 pipelines from the previous version of Milabench and added 3 additional pipelines to address obvious gaps.
+A total of 25 pipelines were suggested, from which 10 were chosen.
+We reused 6 pipelines from the previous version of Milabench and added 3 additional pipelines to address obvious gaps.
 2.3 Selection of benchmarks
@@
 For example, in LLM tasks, pure random noise would result in the LLM consistently producing maximum-length sequences.
-The randomly generated datasets are sized to match popular real-world datasets; for instance, FakeImageNet consists of images sized at (384, 384), based on the average image sizes in ImageNet [4], and FakeVideos consists of 30 FPS, 10-second videos at (640, 480), modeled after the Kinetics dataset [13].
+The randomly generated datasets are sized to match popular real-world datasets; for instance, FakeImageNet consists of images sized at (384, 384), based on the average image sizes in ImageNet [4], and FakeVideos consists of 30 FPS, 10-second videos of size (640, 480), modeled after the Kinetics dataset [13].
 This approach is also applied to model initialization, particularly in inference mode.
@@
 These benchmarks include: MMoCLIP, A multi-modal text-image task using a ViT-L-14 model, trained across 8 nodes for a total of 32 GPUs; Megatron-LM, A 175 billion-parameter language model trained on a language modeling task, trained across 96 nodes for a total of 384 GPUs; and ResNet, A ResNet50 model trained for image classification, executed on 10 nodes for a total of 40 GPUs.
-MMoCLIP and Megatron-LM are excellent candidates to benchmark large-scale training.
-ResNet on the other hand uses a model that is arguably too small to benefit from large scale training on latest GPUs.
-The benchmark is missing important domains of AI such as RL and Graph learning.
-This may be intentional however as the benchmark is tailored for large scale training whereas RL and Graph learning pipelines are more generally executed on a single GPU or a single node.
-While MMoCLIP and Megatron-LM are well-suited to benchmark large-scale AI training, the ResNet benchmark is based on a model that may be too small to fully exploit the capabilities of modern GPUs in large-scale training scenarios.
-Moreover, the benchmark suite does not cover critical AI domains such as reinforcement learning (RL) and graph learning.
+While MMoCLIP and Megatron-LM are excellent candidates to benchmark large-scale AI training, the ResNet benchmark is based on a model that may be too small to fully exploit the capabilities of modern GPUs in large-scale training scenarios.
+The benchmark suite does not cover important AI domains such as reinforcement learning (RL) and graph learning.
 This omission is likely intentional however, as these domains are typically executed on a single GPU or node, whereas the suite is specifically tailored for large-scale training environments.